OpenStudy (anonymous):

Eastside Gym charges a $60 initial fee and $28.50 per month. Valley Gym charges $36 per month, and no initial fee. After how many months of use would Eastside cost less than Valley? A. more than 8 months B. more than 10 months C. less than 7 months D. more than 9 months

OpenStudy (anonymous):

Do you know how to put this question into an equation?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

all right, so, Eastside Gym charges a $60 initial fee and $28.50 per month. This means that f(m) = 60 + 28.5m Valley Gym charges $36 per month, and no initial fee This means that f(m) = 0 + 36m Do you understand?

OpenStudy (anonymous):

(so far) ?

OpenStudy (anonymous):

All I did was putting the question into some functions. Eastside Gym charges a 60 initial fee and 28.5 per month. This means that total cost = 60 + 28.5*number of months. In better notation: f(m) = 60 + 28.5m. Same thing for the other function. Do you understand now?

OpenStudy (anonymous):

so what will be my equation

OpenStudy (anonymous):

Once you do, the next thing to realise is that we need to find the point at which the total cost for eastside gym is equal to the total cost of valley gym. Putting this statement into a mathematical equation, you get: 60+28.5m = 0 + 36m Now solve for m, and m will represent several things: - the point at which the graph of the two lines intersect each other. - the total amount of months, after which the total cost of east gym is equal to the total cost of valley gym. So now you will have the point where they are equal. When will it be less then equal? at anything more than the intersect you just found. You can see this by graphing the two equation, I recommend you to do it if you find this confusing.

OpenStudy (anonymous):

graphing the two functions*

OpenStudy (anonymous):

ok hold on let me solve for m

OpenStudy (anonymous):

will it be 8months

OpenStudy (anonymous):

it will be 60/(36-28.5)

OpenStudy (anonymous):

I think that is 8

OpenStudy (anonymous):

this is my answer choice A. more than 8 months B. more than 10 months C. less than 7 months D. more than 9 months

OpenStudy (anonymous):

Well I am not going to give you the answer. I hope I taught you how to solve these kind of problems though. Can you try find the answer yourself with the explanation I gave you?

OpenStudy (anonymous):

i said its going to be 8 more than 8 months am i right or wrong

OpenStudy (anonymous):

why do you think it is "8 more" than 8 months?

OpenStudy (anonymous):

because i divide 60=7.5

OpenStudy (anonymous):

60/7.5 = ?

OpenStudy (anonymous):

ok then its less than 8

OpenStudy (anonymous):

No. 60/7.5 = 8

OpenStudy (anonymous):

But it is not 8 more than 8 months

OpenStudy (anonymous):

Because at 8 months, we know that the total cost of EG = total cost of VG

OpenStudy (anonymous):

so in fact, anything, even the smallest value, bigger than 8 months, EG will cost less than VG

OpenStudy (anonymous):

so it will be less than then 7

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

im confused than its has to be more then 9

OpenStudy (anonymous):

You realise that EG will cost less than VG at 8 months right? That is what we just solved for!

OpenStudy (anonymous):

thats why i said my answer will be more than 8

OpenStudy (anonymous):

months

OpenStudy (anonymous):

that is correct. But not "8 more than 8". Anything more than 8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem! :) Hope you understand how to solve these problems now! Good luck :)